OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 less than or greater to 7 or 2x greater than or equal to 22 A number line with an open circle on 8, shading to the left, and an open circle on 11, shading to the right. A number line with an open circle on 8, an open circle on 11, and shading in between. A number line with a closed circle on 8, a closed circle on 11, and shading in between. A number line with a closed circle on 8, shading to the left, and a closed circle on 11, shading to the right.

OpenStudy (anonymous):

\[x - 1 \le 7 or 2x \ge 22\]

OpenStudy (anonymous):

\[x \le 6 \]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[x \ge11\]

OpenStudy (anonymous):

sorry my cousin deleted graphs :/